# Black-Box Optimization with Implicit Constraints for Public Policy

### v4

@@ Introduction
The post-decoding process is initiated only for decoder-generated decisions $x$ that are infeasible. Feasible decisions are directly incorporated into the observed feasible set $\mathcal{D}_{f}$. Through iteratively expanding this observed feasible set, we enhance the accuracy of this process by improving the coverage of $\mathcal{D}_{f}$ on the underlying feasible region $\mathcal{S}$. This is further detailed in the ablation study in the appendix.
@@
We provide theoretical analysis for Algorithm 1. Let $\mathtt{enc}:\mathcal{X}\mapsto\mathcal{Z}$ denote the encoder function, $\mathtt{dec}:\mathcal{Z}\mapsto\mathcal{X}$ denote the decoder function, and $g(z)=f(\psi_{\mathcal{D}_{f}}(\mathtt{dec}(z))):\mathcal{Z}\mapsto\mathbbm{R}$ denotes the objective function w.r.t. $\mathcal{Z}$ where $\psi_{\mathcal{D}_{f}}$ is the post-decoder. Note even if post-decoder is not needed, $\psi_{\mathcal{D}_{f}}(\mathtt{dec}(z))=\mathtt{dec}(z)$. Following the existing work in Gaussian process bandit optimization, we utilize cumulative regret and its expected version to evaluate the performance of our algorithm which are defined as follows.
$R_{T}=\sum_{i=1}^{T}f(x_{t})-f(x_{*}),\quad\mathbb{E}[R_{T}]=\mathbb{E}\left[\sum_{i=1}^{T}f(x_{t})-f(x_{*})\right],$ (5)
where $x_{*}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}}f(x)$ and the expectation is taken over all randomness, including random noise and random sampling over observations.
We assume the distance between any two points in $\mathcal{Z}$ can be upper bounded by their distance in $\mathcal{X}$, i.e., $\forall x,x^{\prime}\in\mathcal{X},\|\mathtt{enc}(x)-\mathtt{enc}(x^{\prime})\|_{2}\leq C_{p}\|x-x^{\prime}\|_{2}$. We further assume that function $g:\mathcal{Z}\mapsto\mathbbm{R}$ is drawn from a Gaussian Process prior and it is $C_{g}$-Lipschitz continuous, i.e., $\forall z,z^{\prime}\in\mathcal{Z},|g(z)-g(z^{\prime})|\leq C_{g}\|z-z^{\prime}\|_{2}$. This assumption is the standard Gaussian process assumption for Bayesian optimization (Srinivas et al. 2010). Here $C_{p}$ and $C_{g}$ are universal constants. We use big $\tilde{O}$ notation to omit any constant and logarithmic terms. Now we are ready to state our main theoretical result.
@@ Conclusion
In this paper, we introduce and explore a new category of optimization problems in public policy, termed Implicit-Constrained Black-Box Optimization (ICBBO), where constraints are implicit but can be easily verified. This problem is well motivated by real-world police redistricting problems where districting plans are constrained to socio-economic or political considerations but feasible plans can be easily verified by policymakers. This new ICBBO problem poses more challenges than conventional black-box optimization where existing work cannot be readily applied.
To address this problem, we develop the CageBO algorithm and its core idea is to learn a conditional generative representation of feasible decisions to effectively overcome the complications posed by implicit constraints. In this way, our method can not only learn a good representation of the original space but also generate feasible samples that can be used later. Our theoretical analysis confirms that CageBO can, on expectation, identify the global optimal solution. This approach was empirically validated through a case study on police redistricting in Atlanta, Georgia, where our algorithm demonstrated superior performance in optimizing districting plans. This case study highlights the potential social impact of CageBO, offering a tool that can contribute to more equitable and efficient public policy decisions. Beyond this application, our method is adaptable enough to incorporate traditional constraints by generating compliant samples before optimization, extending its utility across a wide range of policy-making scenarios.
@@ Appendix A Appendix
Function $g:\mathcal{Z}\mapsto\mathbbm{R}$ is drawn from some Gaussian Process and it is $C_{g}$-Lipschitz continuous, i.e.,
@@
Lemma 6 (Expected minimum distance (Lemma 19.2 of Shalev-Shwartz and Ben-David (2014))).
@@
In practice, we approximate the conditional distribution $q(z|x,c)$ and $p(x|z,c)$ by building the ${\mathrm{encoder}}_{\phi}$ and ${\mathrm{decoder}}_{\theta}$ neural networks, and denote the approximated conditional distribution to be $q_{\phi}(z|x,c)$ and $p_{\theta}(x|z,c)$ respectively. In our formulation, the prior of the latent variable is influenced by the feasibility $c$; however, this constraint can be easily relaxed to make the latent variables statistically independent of their labels, adopting a standard Gaussian distribution $\mathcal{N}(0,I)$. Both $q(z|x,c)$ and $p(x|z,c)$ are typically modeled as Gaussian distributions to facilitate closed-form KL divergence computation. Specifically, we introduce generator layers $l(\epsilon,x,c)$ and $l(\epsilon)$ to represent $q(z|x,c)$ and $p(z)$, respectively, transforming the random variable $\epsilon\sim\mathcal{N}(0,I)$ using the reparametrization trick (Sohl-Dickstein et al. 2015). The log-likelihood of the first term is realized as the reconstruction loss, computed with the training data.
@@
It is worth mentioning that this work shares the same concerns with prior research on latent-space Bayesian optimization using VAEs (Notin, Hernández-Lobato, and Gal 2021), where the generated latent variables often fail to consistently represent the underlying data distributions. Such limitations could lead to social impacts similar to those observed with predictive policing systems (Meijer and Wessels 2019), which have faced significant controversy due to their reinforcement of existing biases. Specifically, redistricting to minimize workload variance based on historical 911 calls can inadvertently justify the concentration of policing efforts in districts with a history of higher incident rates. We recognize the critical importance of addressing these ethical concerns, particularly in the context of police redistricting based on historical data, which might inadvertently prioritize areas with historically higher incident rates.
To mitigate these biases, our approach involves debiasing historical data using established machine learning techniques to ensure fairness in districting plans (Zhu and Xie 2022; Jiang and Xie 2024). Furthermore, our CageBO algorithm is designed to accommodate implicit constraints aimed at promoting equitable districting, as suggested by practitioners. These measures align with our commitment to consider the broader impact of our work and the ethical implications of implementing machine learning solutions in public policymaking.

### v5

@@ Introduction
The post-decoding process is initiated only for decoder-generated decisions $x$ that are infeasible. Feasible decisions are directly incorporated into the observed feasible set $\mathcal{D}_{f}$. Through iteratively expanding this observed feasible set, we enhance the accuracy of this process by improving the coverage of $\mathcal{D}_{f}$ on the underlying feasible region $\mathcal{S}$. This is further detailed in the ablation study in the appendix  [^1] Appendix is available on https://arxiv.org/abs/2310.18449.
Code is available on https://github.com/wenqian-xing/CageBO.
@@
We provide theoretical analysis for Algorithm 1. Let $\mathtt{enc}:\mathcal{X}\mapsto\mathcal{Z}$ denote the encoder function, $\mathtt{dec}:\mathcal{Z}\mapsto\mathcal{X}$ denote the decoder function, and $g(z)=f(\psi_{\mathcal{D}_{f}}(\mathtt{dec}(z))):\mathcal{Z}\mapsto\mathbb{R}$ denotes the objective function w.r.t. $\mathcal{Z}$ where $\psi_{\mathcal{D}_{f}}$ is the post-decoder. Note even if post-decoder is not needed, $\psi_{\mathcal{D}_{f}}(\mathtt{dec}(z))=\mathtt{dec}(z)$. Following the existing work in Gaussian process bandit optimization, we utilize cumulative regret and its expected version to evaluate the performance of our algorithm which are defined as follows.
$R_{T}=\sum_{i=1}^{T}f(x_{t})-f(x_{*}),\mathbb{E}[R_{T}]=\sum_{i=1}^{T}\mathbb{E}[f(x_{t})-f(x_{*})],$ (5)
where $x_{*}=\mathop{\mathrm{argmin}}_{x\in\mathcal{X}}f(x)$ and the expectation is taken over all randomness, including random noise and random sampling over observations.
We assume the distance between any two points in $\mathcal{Z}$ can be upper bounded by their distance in $\mathcal{X}$, i.e., $\forall x,x^{\prime}\in\mathcal{X},\|\mathtt{enc}(x)-\mathtt{enc}(x^{\prime})\|_{2}\leq C_{p}\|x-x^{\prime}\|_{2}$. We further assume that function $g:\mathcal{Z}\mapsto\mathbb{R}$ is drawn from a Gaussian Process prior and it is $C_{g}$-Lipschitz continuous, i.e., $\forall z,z^{\prime}\in\mathcal{Z},|g(z)-g(z^{\prime})|\leq C_{g}\|z-z^{\prime}\|_{2}$. This assumption is the standard Gaussian process assumption for Bayesian optimization (Srinivas et al. 2010). Here $C_{p}$ and $C_{g}$ are universal constants. We use big $\tilde{O}$ notation to omit any constant and logarithmic terms. Now we are ready to state our main theoretical result.
@@ Conclusion
In this paper, we introduce a new category of optimization problems in public policy, termed Implicit-Constrained Black-Box Optimization (ICBBO), where constraints are implicit but can be easily verified. This problem is well motivated by real-world police redistricting problems where districting plans are constrained to socio-economic or political considerations but feasible plans can be easily verified by policymakers. This new ICBBO problem poses more challenges than conventional black-box optimization where existing work cannot be readily applied.
To address this problem, we develop the CageBO algorithm and its core idea is to learn a conditional generative representation of feasible decisions to effectively overcome the complications posed by implicit constraints. In this way, our method can not only learn a good representation of the original space but also generate feasible samples that can be used later. Our theoretical analysis confirms that CageBO can, on expectation, identify the global optimal solution. This approach was empirically validated through a case study on police redistricting in Atlanta, Georgia, where our algorithm demonstrated superior performance in optimizing districting plans.
Acknowledgments
This work was partially done when CL was at the University of Chicago. This work was partially supported by the UAlbany Computer Science Department startup funding. The authors would like to thank the AAAI-AISI reviewers and area chair for helpful comments.
@@ Appendix A Appendix
Function $g:\mathcal{Z}\mapsto\mathbb{R}$ is drawn from some Gaussian Process and it is $C_{g}$-Lipschitz continuous, i.e.,
@@
Lemma 6 (Expected minimum distance (Lemma 19.2 of shalev2014understanding)).
@@
In practice, we approximate the conditional distribution $q(z|x,c)$ and $p(x|z,c)$ by building the ${\mathrm{encoder}}_{\phi}$ and ${\mathrm{decoder}}_{\theta}$ neural networks, and denote the approximated conditional distribution to be $q_{\phi}(z|x,c)$ and $p_{\theta}(x|z,c)$ respectively. In our formulation, the prior of the latent variable is influenced by the feasibility $c$; however, this constraint can be easily relaxed to make the latent variables statistically independent of their labels, adopting a standard Gaussian distribution $\mathcal{N}(0,I)$. Both $q(z|x,c)$ and $p(x|z,c)$ are typically modeled as Gaussian distributions to facilitate closed-form KL divergence computation. Specifically, we introduce generator layers $l(\epsilon,x,c)$ and $l(\epsilon)$ to represent $q(z|x,c)$ and $p(z)$, respectively, transforming the random variable $\epsilon\sim\mathcal{N}(0,I)$ using the reparametrization trick (sohl2015deep). The log-likelihood of the first term is realized as the reconstruction loss, computed with the training data.
@@
It is worth mentioning that this work shares the same concerns with prior research on latent-space Bayesian optimization using VAEs (notin2021improving), where the generated latent variables often fail to consistently represent the underlying data distributions. Such limitations could lead to social impacts similar to those observed with predictive policing systems (meijer2019predictive), which have faced significant controversy due to their reinforcement of existing biases. Specifically, redistricting to minimize workload variance based on historical 911 calls can inadvertently justify the concentration of policing efforts in districts with a history of higher incident rates. We recognize the critical importance of addressing these ethical concerns, particularly in the context of police redistricting based on historical data, which might inadvertently prioritize areas with historically higher incident rates.
To mitigate these biases, our approach involves debiasing historical data using established machine learning techniques to ensure fairness in districting plans (Zhu and Xie 2022; jiang2024graph). Furthermore, our CageBO algorithm is designed to accommodate implicit constraints aimed at promoting equitable districting, as suggested by practitioners. These measures align with our commitment to consider the broader impact of our work and the ethical implications of implementing machine learning solutions in public policymaking.